OpenStudy (anonymous):

Calculation of time: How much time is required to produce 6,500 C of charges using a current of 15 A? I know that A = C/s , but I'm not quite sure how to go about solving and setting up the actual equation. Thank you for your help!

OpenStudy (anonymous):

s = C /A = 6500 C / 15 A = 6500 C / (15 C/s) = 433.33 s or 7.22 min

OpenStudy (anonymous):

Oh! That makes sense now that I see it out. Thank you :)

OpenStudy (anonymous):

No problem (: